A boy is x inches tall now, if his height increased by 10% a year. Which of the following expressions represents his height three years from now?( answer is 1.331x show your solution)
Answer by Alan3354(69443)

Nothing followed.
--------------------
h(t) = h*(1 + 0.1)^t (t in years)
h(t) = h*1.1^3
= 1.331*h